OpenStudy (anonymous):

divide 60 kg in the ratio of 5 : 7.

OpenStudy (anonymous):

60/12=5 so the shares are 5*5=25 5*7=35

OpenStudy (anonymous):

what does the * mean . I am lost on the shares past

OpenStudy (anonymous):

Is the answer 25:35?